If a solution of nacl contains 12.5 grams per 100 grams of solution and has density of 1.09 grams per cms.express it concentration in molarity, per cent in weight,morality,mole per cent
1
Expert's answer
2019-10-07T06:43:22-0400

1)Molarity

"C=\frac{m_{NaCl}·\rho ·1000 }{M_{NaCl}·m} =\frac{12.5·1.08·1000}{58.5·112.5}=2.0M"

2) weight %

"\omega _{NaCl}=\frac{m_{NaCl}}{m_{H_2O}+m{NaCl}}·100=\frac{12.5}{112.5}·100=11.11" %

3) Molality

"C_m=\frac{m_{NaCl}·}{M_{NaCl}·m_{H_2O(kg)}}=\frac{12.5}{58.5·0.1}=2.1m"

4)mol %

"\chi=\frac{\frac{m_{NaCl}}{M_{NaCl}}}{\frac{m_{NaCl}}{M_{NaCl}}+\frac{m_{H_2O}}{M_{H_2O}}}·100=\frac{0.214}{0.214+5.556}·10=3.7"%